### v1

# Rational Design Heterobilayers Photocatalysts for Efficient Water Splitting Based on 2D Transition-Metal Dichalcogenide and Their Janus

Abstract
Direct Z-scheme heterostructures with enhanced redox potential are increasingly regarded as promising materials for solar-driven water splitting. This potential arises from the synergistic interaction between the intrinsic dipoles in Janus materials and the interfacial electric fields across the layers. In this study, we explore the photocatalytic potential of 20 two-dimensional (2D) Janus transition metal dichalcogenide (TMDC) heterobilayers for efficient water splitting. Utilizing density functional theory (DFT) calculations, we first screen these materials based on key properties such as band gaps and the magnitude of intrinsic electric fields to identify promising candidates. We then evaluate additional critical factors, including carrier mobility and surface chemical reactions, to fully assess their performance. The intrinsic dipole moments in Janus materials generate built-in electric fields that enhance charge separation and reduce carrier recombination, thereby improving photocatalytic efficiency. Furthermore, we employ the Fröhlich interaction model to quantify the mobility contributions from the longitudinal optical phonon mode, providing detailed insights into how carrier mobility, influenced by phonon scattering, affects photocatalytic performance. Our results reveal that several Janus-TMDC heterobilayers, including WSe2-SWSe, WSe2-TeWSe, and WS2-SMoSe, exhibit strong absorption in the visible spectrum and achieve solar-to-hydrogen (STH) conversion efficiencies of up to 33.24%. These findings demonstrate the potential of Janus-based Z-scheme systems to overcome existing limitations in photocatalytic water splitting by optimizing the electronic and structural properties of 2D materials. This research highlights a viable pathway for advancing clean energy generation through enhanced photocatalytic processes.
keywords:
Janus TMDC, photocatalyst, water-spliting, carrier mobility, Fröhlich interaction
[…]
1 Introduction
Due to escalating environmental pollution and energy challenges, the search for alternative energy sources has gained significant attention in the modern era. Hydrogen energy is increasingly regarded as an innovative technology for the future, providing the dual advantages of being both clean and renewable. 1, 2, 3. Photocatalytic water splitting has garnered significant attention in global scientific and industrial communities, which convert solar energy through the splitting process of water to H2 energy. This utilization of solar light energy is accelerating the carbon neutralization process in the world 4. The photocatalytic technique is based on generating charge carriers within a photocatalyst system under suitable light irradiation. Therefore, the primary challenge is discovering a highly efficient photocatalyst that exhibits favorable photoelectronic and optical properties for STH energy conversion. 5, 6. Maximizing the efficiency of photocatalytic hydrogen generation depends on the control optimization of the critical parameters, including i) an optimal band gap and appropriately aligned conduction and valence band of the photocatalyst to effectively harvest visible light and facilitate the redox reactions required for water splitting; ii) a high surface area and well-engineered morphology, alongside a crystalline structure, enhancing light absorption and efficiently promoting charge mobility; iii) suppression of electron-hole recombination and enhanced charge carrier mobility for efficient charge transport to reactive sites; iv) reaction conditions, including pH, temperature, and the incorporation of sacrificial agents 6, 5, 7, 8.
Among the various photocatalysts, two-dimensional (2D) materials offer many favorable properties, such as large surface area, quantum confinement effects, strong light absorption enhanced charge separation, and structural flexibility 9, 10, 11, 12. One of the standout features of 2D materials is their planar structure, which offers a high density of surface atoms and ample space for integration with co-catalysts. This results in more available active sites for catalytic reactions, and the potential creates new active sites, enhancing overall catalytic efficiency 13. Additionally, the significantly reduced thickness of 2D materials enables charge carriers to travel shorter distances to reach the surface. This leads to decreasing electron-hole recombination and preserves more charge carriers for surface reactions. These unique characteristics make 2D materials promising candidates for improving the performance of photocatalytic systems 13, in which TMDC materials have been considered as the most potential applicant due to possessing efficient visible light absorption and high charge carrier mobility enhancing charge separation. The TMDCs have demonstrated good chemical stability and robust light-matter interactions, further improving their efficacy in light-driven processes. Despite these strengths, TMDCs are facing plausible limitations in photocatalytic hydrogen generation due to high exciton recombination and strong electrostatic coulomb interactions, leading to inefficient photogenerated charge carriers 14, 15. In comparison with conventional TMDCs, Janus TMDCs include different chalcogen atoms, breaking the out-of-plane symmetry 16. Thus, the Janus TMDCs will create out-of-plane dipoles between the top and bottom layers that could generate an intrinsic electric field and enhance the photocatalytic performance in Janus-TMDCs 17, 18, 19.
Based on the above reasons, the heterobilayers combining TMDCs and their Janus materials that have been successfully synthesized 20, 21 could benefit photocatalytic water splitting. In monolayer 2D material, both the hydrogen evolution reaction (HER) and the oxidation evolution reaction (OER) occur on the same surface of the 2D material, resulting in a higher recombination rate of photogenerated charge carriers 6, 5. A promising solution to this problem is the van der Waals (vdW) heterobilayers of two semiconductor layers. In particular, the valence band maximum (VBM) and conduction band minimum (CBM) are integrated into both semiconductor sides, which are known as type II and Z-scheme heterostructure systems. A distinct behavior of the Z-scheme and type-II heterostructures is their unique electronic structure. Notably, the Z-scheme junction features a suitable band alignment and a wide directional built-in electric field. As a result, photogenerated charge carriers with a large redox potential are well-preserved and spatially separated within the Z-scheme heterostructures. Combining two narrow band gap semiconductors into a Z-scheme system could solve the challenge of a single platform or type-II junctions, where strong redox ability and light absorption region may not be obtained simultaneously 22, 23. This configuration allows HER and OER activity on separate surfaces. On the other hand, the heterobilayer with the Janus structure generates an intrinsic electric field, which not only enhances the photocatalytic performance but also accelerates the reaction kinetics by forming synergistic surface properties that allow selective adsorption of water molecules on each surface 24, 25, 26, 27. Moreover, the configuration of the heterobilayer can determine whether photogenerated carriers follow a Z-scheme or a traditional type-II pathway 27. However, the diversity of possible Janus-TMDC combinations in heterobilayers poses a significant challenge in determining the optimal configuration for photocatalytic water splitting. In this study, we provide a theoretical framework for rapidly identifying effective Janus-TMDC pairs through a top-down approach: candidates are initially filtered based on band alignment, and further refinement follows as needed. Another key challenge is the intrinsic polarization of Janus materials, which uniquely affects their electrical mobility. However, this factor is often neglected in previous Janus or TMDC studies for photocatalytic 28, 1, 29. To address this, we have adopted Fröhlich interaction model 30, 28 to account for the mobility adjustments in Janus heterostructures accurately.
[…]
2 Theoretical Approach and Computational Details
In this research, we employed DFT for calculations using the Quantum ESPRESSO package 31, 32, 33. The generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional was utilized as implemented in the standard solid-state pseudopotentials (SSSP) library 34, 35, 36. Van der Waals interactions were incorporated using the Grimme DFT-D2 method 37, 38, 39. To address the underestimation by PBE in predicting band edge positions and band gaps, we used the Heyd-Scuseria-Ernzerhof (HSE) hybrid-functional with a mixing parameter 25% 40, 41. The kinetic cutoff energy for wavefunctions ($E_{\text{cut}}$) was set to 800 eV, and the Monkhorst-Pack k-point grid was set to $12\times 12\times 1$ for PBE and $6\times 6\times 1$ for HSE based on the convergence test. All structures were fully relaxed using the BFGS scheme 42, with convergence criteria for force and energy differences set to $1\times 10^{-4}$ eV/Å and $1\times 10^{-8}$ eV, respectively. To prevent interlayer interactions, a relatively large vacuum layer of 30 Å was considered along the $z$-direction. To examine the Gibbs free energy and the mobility of the hydrogen atom, we adopted a $3\times 3\times 1$ supercell with 54 atoms. For the carrier mobility calculation, a rectangle $\sqrt{3}\times 1\times 1$ supercell was used to obtain the effective mass of carriers along $x$- and $y$-directions. The Wannier90 code and the VESTA software were used for pre-processing and post-processing 43, 44.
3 Results and Discussion
3.1 Geometric and Electronic Structures
The Janus-TMDC heterobilayers have many configurations by selecting the atoms at the interface between two layers. Replacement of these atoms results in changes in the difference in electronegativity between layers, which, in turn, modifies their optoelectrical properties and influences the photocatalytic capabilities of the material. Based on this concept, Janus MoSSe, WSSe, and their heterobilayers with TMDC, such as MoSSe-MoS2, have already been successfully synthesized 20, 21. Following these achievements, recent theoretical research has focused on other Janus-type materials based on molybdenum (Mo) and tungsten (W) 45, 46, 47, 48, 49, 50. Thus, we selected Mo and W as the metals of interest because of their promising possibilities for future synthesis and applications in water splitting. For the chalcogenides, we used combinations of sulfur (S), selenium (Se), and tellurium (Te), resulting in the formation of 2D vdW structures comprising conventional TMDC and Janus materials.
Figure 1: Visual representations of the vdW Janus-TMDC structure and its Brillouin zone: (a) top view; (b) side view; (c) Brillouin zone indicating high-symmetry points.
The geometric structures of these 20 materials were designed based on the previously synthesized AA stacking mode of the MoSSe-MoS2 structure, systematically replacing the metal and chalcogenide atoms, as shown in Figure 1. The optimized lattice parameters are listed in Table 1. We note that the structures of MoS2-TeMoSe and WS2-TeWSe did not converge during the optimization calculations, indicating that these configurations are unstable. This instability can be attributed to the significant differences in electronegativity between the S and Te atoms.
3.2 Photocatalytic characteristic
After completion of structural optimizations, we evaluated the potential of 18 materials as photocatalysts (see Table S1 in Supporting Information). To determine the electronic properties critical for photocatalytic activity, we performed partial energy band structure for 18 material combinations, in which the partial band structures show the contributions of individual layers in the heterostructures to the electronic states near the band edges (see Figure S1, Supporting Information). By projecting the electronic states onto the atomic orbitals of the Janus and TMDC layers, we identified which layers predominantly contribute to the VBM and CBM. This approach allowed us to classify the materials into Type I band alignment, where both the VBM and CBM are contributed by the same layer, or Type II alignment and Z-scheme alignment, where they are contributed by different layers, based on the spatial localization of their VBM and CBM.
[…]
To be considered as a photocatalyst for water splitting, the theoretical band gap should be larger than 1.23 eV to overcome the reduction and oxidation potential of H2O 1, 51. However, due to the difference in electrostatic potential between the layers contributing to the VBM and CBM, even materials with smaller band gaps can act as photocatalysts. As shown in Table S1 (Supporting Information), the materials with Type I band alignment, including MoSe2-SeMoTe, WSe2-SeWS, and WSe2-SeWTe, have band gaps less than 1.23 eV. Thus, they are unsuitable for water-splitting applications.
Figure 2: HSE band structure and band alignment at pH = 7 for selected Janus-TMDC heterobilayers: WSe2-SWSe, WSe2-TeWSe, and WS2-SMoSe.
For the remaining materials, we calculated the electrostatic potential curves, as shown in Figure S2 in Supporting Information. The calculations include a dipole correction to suppress artificial dipole fields by introducing an additional ramp-shaped potential 42. The intrinsic electric field generated by the Janus layer in the heterostructure results in a significant potential difference between the two layers, leading to a difference in work function ($\Delta\Phi$) as calculated and presented in Table 1. Note that $\Delta\Phi$ is determined by taking the difference between the work functions of the TMDC and Janus layers, following this equation 14:
$\Delta\Phi=V_{\text{TMDC}}-V_{\text{Janus}}.$ (1)
The band alignment plays a critical role in determining the photocatalytic properties of Janus-TMDC heterobilayers. When the work function difference between the layers, denoted as $\Delta\Phi$, is positive ($\Delta\Phi>0$), this indicates that the vacuum level of the TMDC layer is higher than that of the Janus layer. In such a configuration, the TMDC layer predominantly contributes to the VBM, while the Janus layer largely influences the CBM. This arrangement leads to a Z-scheme band alignment, which effectively broadens the separation between the CBM and VBM, enhancing charge carrier separation and improving photocatalytic activity (e.g., WSe2-SWSe and WS2-SMoSe, as shown in Figure 2). Conversely, when $\Delta\Phi$ is negative ($\Delta\Phi<0$), meaning the vacuum level of the Janus layer is higher than that of the TMDC layer, the Janus layer mainly contributes to the VBM, and the TMDC layer to the CBM. This configuration also results in a Z-scheme alignment (e.g., WSe2-TeWSe, as shown in Figure 2).
In cases where $\Delta\Phi$ is negligible or does not meet the conditions above, the system exhibits a type II band alignment. The effective band gap in a Z-scheme configuration can be expressed as $E_{\text{eff}}=E_{g}+\Delta\Phi$, where $E_{g}$ represents the intrinsic band gap, and $\Delta\Phi$ reflects the built-in potential difference at the heterostructure interface. Materials with $E_{\text{eff}}>1.23$ eV are considered promising candidates for photocatalytic applications, and their performance is further evaluated through calculations of STH efficiency. The $E_{\text{eff}}$ values for the materials studied are summarized in Table 1.
Table 1: Calculated band gap energies, work function differences, and effective band gaps for various Janus-TMDC heterobilayers.
[…]
MoS2-SeMoTe 0.315 0.82 II -
MoS2-TeMoSe - - - -
MoSe2-SeMoTe 1.044 0.84 I 1.044
MoSe2-TeMoSe 0.481 -0.6 Z-scheme 1.081
MoS2-SMoSe 1.313 0.82 I 1.313
MoS2-SeMoS 0.68 -0.64 Z-scheme 1.32
MoSe2-SMoSe 0.494 0.71 Z-scheme 1.204
MoSe2-SeMoS $1$0.73 I 1.111
WS2-SWSe 1.437 0.79 I 1.437
WS2-SeWS 0.815 -0.66 Z-scheme 1.475
WSe2-SWSe 0.604 0.69 Z-scheme 1.294
WSe2-SeWS 1.211 -0.7 I 1.211
WS2-SeWTe 0.411 0.93 II -
WS2-TeWSe - - - -
WSe2-SeWTe 1.103 0.8 I 1.103
WSe2-TeWSe 0.969 -0.63 Z-scheme 1.599
MoS2-SWSe 1.562 0.83 II 0.732
MoS2-SeWS 0.952 -0.58 Z-scheme 1.532
WS2-SMoSe 0.937 0.79 Z-scheme 1.717
WS2-SeMoS 1.517 -0.71 Z-scheme 2.227
3.3 Solar to Hydrogen Conversion Efficiency (STH)
After screening all possible materials for their ability to initialize water splitting, we identified 9 materials (see Table 2) with an effective band gap greater than 1.23 eV, indicating that these materials can potentially straddle the whole redox potential of H2O. The redox potential of water depends on pH and can be described as follows:
$E_{\text{H}^{+}/\text{H}_{2}}=0-0.059\times\text{pH},$ (2)
and
$E_{\text{O}_{2}/\text{H}_{2}\text{O}}=1.23-0.059\times\text{pH}.$ (3)
Determining the appropriate pH for water splitting is crucial for optimizing the performance of photocatalysts. Therefore, we analyzed the band alignment of the selected materials for the redox potentials at $\text{pH}=0$ and $7$, as illustrated in Figure S3 (Supporting Information). We find that some materials require solution $\text{pH}<-2$, indicating highly acidic conditions, which limits their practical application 52, 53. Therefore, these materials are excluded from the STH efficiency calculations. Only materials capable of splitting water at $\text{pH}=0$ and $7$ are considered for the STH efficiency evaluation, with the relevant material information provided in Table 2. It is noted that the VBM of WSe2-SWSe is slightly higher than the OER potential. Therefore, at $\text{pH}=7$, this material exhibits only the hydrogen evolution half-reaction during photocatalysis.
Table 2: STH conversion efficiencies, including overpotentials $(\chi_{\text{H}}$ and $\chi_{\text{O}})$, absorption $(\eta_{\text{abs}})$ and carrier utilization efficiencies $(\eta_{\text{cu}})$ for selected materials.
Material $\chi_{\text{H}}$ $\chi_{\text{O}}$ $\eta_{\text{abs}}$ $\eta_{\text{cu}}$ $\eta_{\text{STH}}$
WSe2-SWSe 0.1 0.01 96.63% 45.50% 30.71%
WSe2-TeWSe 0.12 0.25 87.08% 46.45% 30.72%
WS2-SMoSe 0.09 0.39 88.54% 52.77% 33.24%
WS2-SeMoS 0.27 0.73 60.68% 57.62% 29.12%
MoS2-SMoSe Require low pH (pH $<$ 0)
MoS2-SeMoS
WS2-SWSe
WS2-SeWS
MoS2-SeWS
The STH efficiency $\eta_{\text{STH}}$ of the selected materials were calculated using the following expression 1:
$\eta_{\mathrm{STH}}=\eta_{\mathrm{abs}}\times\eta_{\mathrm{cu}}\times\frac{\left.\int_{0}^{\infty}P(h\omega)\mathrm{d}(h\omega)\right)}{\int_{0}^{\infty}P(h\omega)\mathrm{d}(h\omega)+\Delta\Phi\int_{E_{\mathrm{g}}}^{\infty}\frac{P(h\omega)}{h\omega}\mathrm{d}(h\omega)},$ (4)
where $\eta_{\text{abs}}$ is the light absorption efficiency, $\eta_{\text{cu}}$ is the carrier utilization efficiency, $E_{g}$ is the band gap energy, $\Delta\Phi$ is the electric potential difference, and $P(h\omega)$ represents the AM1.5G solar flux at photon energy $h\omega$. The light absorption efficiency ($\eta_{\text{abs}}$) is given by
$\eta_{\mathrm{abs}}=\frac{\int_{E_{\mathrm{g}}}^{\infty}P(h\omega)\mathrm{d}(h\omega)}{\int_{0}^{\infty}P(h\omega)\mathrm{d}(h\omega)},$ (5)
and the carrier utilization $\eta_{\text{cu}}$ is determined by
$\eta_{\mathrm{cu}}=\frac{\Delta E_{\text{spliting}}\int_{E_{\min}}^{\infty}\frac{P(h\omega)}{h\omega}\mathrm{d}(h\omega)}{\int_{E_{\mathrm{g}}}^{\infty}P(h\omega)\mathrm{d}(h\omega)}$ (6)
where $\Delta E_{\text{spliting}}$ = 1.23 eV for water splitting and
$E_{\text{min }}=\left\{\begin{array}[]{cc}\mathrm{E}_{\mathrm{g}}&\text{ for }\chi\left(\mathrm{H}_{2}\right)\geq 0.2,\chi\left(\mathrm{O}_{2}\right)\geq 0.6\\ E_{\mathrm{g}}+0.2-\chi\left(\mathrm{H}_{2}\right)&\text{ for }\chi\left(\mathrm{H}_{2}\right)<0.2,\chi\left(\mathrm{O}_{2}\right)\geq 0.6\\ E_{\mathrm{g}}+0.6-\chi\left(\mathrm{O}_{2}\right)&\text{ for }\chi\left(\mathrm{H}_{2}\right)\geq 0.2,\chi\left(\mathrm{O}_{2}\right)<0.6\\ E_{\mathrm{g}}+0.8-\chi\left(\mathrm{H}_{2}\right)-\chi\left(\mathrm{O}_{2}\right)&\text{ for }\chi\left(\mathrm{H}_{2}\right)<0.2,\chi\left(\mathrm{O}_{2}\right)<0.6\end{array}\right.,$ (7)
where $\chi(\text{H}_{2})$ and $\chi(\text{O}_{2})$ are the overpotentials for the hydrogen and oxygen evolution reactions, respectively. It is noted that $\chi(\text{H}_{2})$ is the potential difference between the CBM and the redox potential of $\text{H}^{+}/\text{H}_{2}$, while $\chi(\text{O}_{2})$ is the potential difference between the VBM and the redox potential of $_{2}/\text{H}_{2}\text{O}$ O, as shown in Figure 2. Our calculated results, presented in Table 2, show that the selected materials exhibit remarkable photocatalytic efficiencies and exceed the conventional theoretical limit of 18% 54. In particular, WS2-SMoSe achieves an $\eta_{\text{STH}}$ of 33.24%. In comparison, materials reported in previous studies, such as Penta-PdSeTe (29.71%) 55, Janus $\gamma$-GeSSe (28.78%) 1, and SPtGeAs2 (30.74%) 56, show competitive but lower efficiencies.
The exceptional performance of WS2-SMoSe and other materials in Table 2 can be attributed to its optimal electronic structure and favorable band alignment, which facilitate efficient light absorption and charge carrier utilization. The light absorption efficiency $\eta_{\text{abs}}$ is particularly high for materials with smaller band gaps, as shown in Equation 5. WSe2-SWSe with a band gap of 0.604 eV exhibits an outstanding $\eta_{\text{abs}}$ of 96.63%, which allows it to absorb a broader spectrum of solar radiation compared to traditional photocatalysts with a standard band gap of 1.23 eV. It is important to note that Equation 5 assumes the absorbance as a step function zero for photon energies below the band gap and unity for energies above it. This simplification may lead to underestimating $\eta_{\text{abs}}$, particularly at lower photon energies. Therefore, while the calculated $\eta_{\text{abs}}$ provides valuable insights, it may be more accurate for materials with high photon energy absorption.
3.4 Carrier mobility
The relationship between carrier mobility and photocatalytic efficiency is essential in determining the performance of photocatalysts. After the electron-hole pairs are separated, the free photoexcited carriers need to migrate to the chemical active sites for water splitting. Higher mobility enables these charge carriers to travel faster and further, reducing the probability of recombination before they can reach the reactive surface sites 57, 58, 59, 60. Mobility is primarily limited by phonon scattering, which impedes the movement of charge carriers. Traditional approaches to calculating mobility, such as deformation potential theory (DPT), focus on the scattering caused by longitudinal acoustic (LA) phonons 14, 60, 56. However, recent studies, such as Cheng and Liu have demonstrated that this method often underestimates the influence of longitudinal optical (LO) phonon scattering, particularly in Janus materials with strong polar effects.
The carrier mobility $\mu_{\text{LA}}$ due to the LA phonon scattering within the DPT theory is given as follows 59, 1:
$\mu_{\text{LA}}=\frac{e\hbar^{3}C_{2\mathrm{D}}}{k_{\mathrm{B}}Tm^{*}_{x(y)}m_{\mathrm{d}}E_{\mathrm{d}}^{2}},$ (8)
where the effective masses along the zigzag ($x$) and armchair ($y$) directions are represented by $m_{x}^{*}$ and $m_{y}^{*}$, respectively, and the average effective mass given by $m_{d}=\sqrt{m_{x}^{*}m_{y}^{*}}$. A $\sqrt{3}\times 1$ supercell (see Figure S4 in Supporting Information) is used to compute $m_{x}^{*}$ and $m_{y}^{*}$. The in-plane Young’s modulus, $C_{2\mathrm{D}}$, is defined as 59, 1:
$C_{2\mathrm{D}}=\frac{\partial^{2}E_{\text{total}}}{\partial\varepsilon^{2}}/S_{0},$ (9)
where $E_{\text{total}}$ is the total energy after applying a uniaxial strain (with strain $\varepsilon=\Delta l/l_{0}$), and $S_{0}$ is the equilibrium area of the system (see Figure S5, Supporting Information). The deformation potential constant, $E_{d}$, is given by
$E_{d}=\frac{\Delta E}{\varepsilon},$ (10)
where $\Delta E$ is the shift in the energy level of either the CBM or the VBM relative to the vacuum level (see Figure S6 in Supporting Information). The calculated parameters at room temperature ($T=300$ K) are presented in Table 3.
Table 3: Material properties of selected Janus-TMDC heterobilayers, including effective masses ($m^{*}$), deformation potentials ($E_{d})$, and calculated carrier mobilities ($\mu_{\text{LA}}$).
Sample Direction Carrier $m*/m_{0}$ $E_{d}$ $C_{\text{2D}}$ $\mu_{\text{LA}}$
[…]
WSe2-SWSe zigzag h 0.401 2.26 218.7755 5085.70
e 0.225 2.70 11120.74
armchair h 0.499 2.18 201.1680 4038.86
e 0.290 2.64 8298.47
WSe2-TeWSe zigzag h 0.611 2.52 212.9836 1735.10
e 0.257 2.16 13274.89
armchair h 0.743 2.46 199.0206 1399.13
e 0.316 2.26 9215.51
WS2-SMoSe zigzag h 0.252 4.70 245.0984 3499.11
e 0.343 4.70 1697.92
armchair h 0.285 4.78 229.1263 2796.32
e 0.480 4.34 1330.22
[…]
Table 4: Calculated mobilities contributed by LO1 and LO2 modes ($\mu_{\text{LO1}}$ and $\mu_{\text{LO2}}$) and combined LA and LO modes ($\mu_{\text{total}}$).
Sample Direction Carrier $\mu_{\text{LO1}}$ $\mu_{\text{LO2}}$ $\mu_{\text{total}}$
[…]
WSe2-SWSe zigzag h 980.78 5669.29 718.07
e 1719.25 9937.93 1295.01
armchair h 980.78 5669.29 692.72
e 1719.25 9937.93 1245.67
WSe2-TeWSe zigzag h 946.32 246.97 175.99
e 2237.96 584.06 447.56
armchair h 946.32 246.97 171.81
e 2237.96 584.06 441.01
WS2-SMoSe zigzag h 109174.17 2137.68 1311.06
e 72065.71 1411.08 762.48
armchair h 109174.17 2137.68 1198.22
e 72065.71 1411.08 678.28
It is important to note that many 2D TMDCs exhibit mobility limited by not only LA but also LO phonon scattering 28. The LO phonon scattering is closely linked to the Born effective charge, which reflects the degree of electric polarization change due to atomic vibrations. Thus, a large Born effective charge correlates with stronger carrier scattering, leading to reduced mobility. This critical insight highlights the need to go beyond DPT to accurately assess the intrinsic mobility of Janus-TMDC heterostructures when LO phonon scattering could play a substantial role. In the present study, the Fröhlich interaction model 28, 30 is employed to quantify the mobility contributions from the LO phonon mode. The total scattering rate near the band edges can be written as
$\begin{split}\frac{1}{\tau_{\text{LO}}}&=\frac{e^{2}n_{h\omega}}{M_{M}M_{X1}M_{X2}S_{0}d_{eff}^{2}(\varepsilon\varepsilon_{0})^{2}(\hbar\omega)^{2}}\times\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left[\left|Z_{MB}\right|\sqrt{M_{X1}M_{X2}}+\left|Z_{X1}\right|\sqrt{M_{M}M_{X2}}+\left|Z_{X2}\right|\sqrt{M_{M}M_{X1}}\right]^{2},\end{split}$ (11)
where $e$ is the elementary charge, $n_{h\omega}$ is the Bose-Einstein distribution at the phonon frequency at the gamma point, $A$ is the equilibrium area, $d_{eff}$ is the effective thickness, $\varepsilon$ is the in-plane optical dielectric constant, and $\varepsilon_{0}$ is the dielectric constant. Additionally, $M_{M}$, $M_{X1}$, and $M_{X2}$ are the atomic masses of the metal and the two chalcogens in the structure, while $Z_{MB}$, $Z_{X1}$, and $Z_{X2}$ are the Born effective charges of the metal and the two chalcogen atoms (see Table S2 in Supporting Information for reference). Thus, the mobility contributed by the LO phonon mode is given by
$\mu_{\text{LO}}={|e|\tau_{\text{LO}}}/{m^{*}}.$ (12)
In Janus-TMDC heterobilayers, two distinct LO phonon modes, LO1 and LO2, correspond to the vibrations of the TMDC and Janus layers, respectively, as shown in Figure 3. The calculated mobilities for $\mu_{\text{LO1}}$ and $\mu_{\text{LO2}}$ are provided in Table 4. Then the total mobility $\mu_{\text{total}}$ is then determined using Matthiessen’s rule as follows 28:
$\frac{1}{\mu_{\text{total}}}=\frac{1}{\mu_{\text{LO1}}}+\frac{1}{\mu_{\text{LO2}}}+\frac{1}{\mu_{\text{LA}}}.$ (13)
The calculated $\mu_{\text{LO1}}$, $\mu_{\text{LO2}}$, and $\mu_{\text{total}}$ are listed in Table  4. Our results reveal that LO phonon scattering significantly limits mobility in Janus-TMDC heterobilayers, which leads to lower carrier mobility compared to those predicted by DPT alone (see Table  3).
3.5 Optical properties
Figure 4: Optical absorption spectra along in-plane ($E_{\parallel}$) and out-of-plane ($E_{\perp}$) of WSe2-SWSe, WSe2-TeWSe, and WS2-SMoSe. The dashed line denotes the range of the visible light.
The optical property is another factor in determining the effectiveness of semiconductor materials as optoelectronic devices. Additionally, photocatalysts with strong light absorption capabilities will perform more efficiently.
The optical properties of $\mathrm{WSe_{2}}$-SWSe, $\mathrm{WSe_{2}}$-TeWSe, and $\mathrm{WS_{2}}$-SMoSe are calculated by using the independent particle approximation. The absorption coefficient $\alpha(\omega)$ is obtained from the calculated real $\varepsilon_{1}$ and imaginary $\varepsilon_{2}$ parts of the dielectric function as follows 42:
$\alpha(\omega)=\frac{\sqrt{2}\omega}{c}\left\{\sqrt{\varepsilon_{1}^{2}(\omega)+\varepsilon_{2}^{2}(\omega)}-\varepsilon_{1}(\omega)\right\}^{1/2}.$ (14)
Here, the values of $\varepsilon_{1}$ and $\varepsilon_{2}$ are scaled by the factor $h/d_{0}$ where $h$ represents the unit cell thickness and $d_{0}$ denotes the effective material thickness 14. The effective thickness, $d_{0}$, is calculated using the formula: $d_{0}=d_{\text{outer}}+r_{\text{atom1}}+r_{\text{atom2}}$, where $d_{\text{outer}}$ is the distance between the two outermost atoms of the heterostructure, as illustrated in 1. The terms $r_{\text{atom1}}$ and $r_{\text{atom2}}$ correspond to the vdW radii of the two outermost atoms 61.
The absorption spectra exhibit strong anisotropic behavior, with $\alpha(\omega)$ along the in-plane ($E_{\parallel}$) reaching up to $0.84\times 10^{6}$ cm-1 in the visible light region for $\mathrm{WS_{2}}$-SMoSe and WSe2-TeWSe, and $0.77\times 10^{6}$ cm-1 for WSe2-SWSe. These values surpass those of monolayer $\gamma$-GeSSe ($0.44\times 10^{6}$ cm-1) 1 and PdSSe ($0.24\times 10^{6}$ cm-1) 62. Notably, the $E_{\parallel}$ optical absorption differs considerably from the out-of-plane ($E_{\perp}$) absorption, especially within the visible spectrum. The combination of superior STH efficiency and strong light absorption across all materials highlights their potential for optoelectronic and water-splitting applications.
3.6 Hydrogen Evolution Reaction
3.6.1 Gibbs free energy steps for HER
Figure 5: Calculated Gibbs free energy steps for HER under various illumination conditions.
The research indicates that the selected photocatalysts exhibit characteristics of direct Z-scheme systems, featuring appropriate band edge potentials and high efficiency. However, these properties alone do not ensure the spontaneous occurrence of the water-splitting reaction under light irradiation. To further evaluate the hydrogen evolution reaction (HER) capabilities, the Gibbs free energy change ($\Delta G_{\text{H}}$) for hydrogen adsorption is considered a crucial factor in assessing the HER capabilities of photocatalysts. To accurately calculate the adsorption energy and minimize interactions between adjacent atoms, a 3 $\times$ 3 $\times$ 1 supercell was constructed for each material, ensuring that the results reflect the behavior of isolated hydrogen atoms. The Gibbs free energy change was determined using the equations:
$\Delta G_{\text{H}}=\Delta E_{\text{H}}+0.24\text{ eV},$ (15)
and
$\Delta E_{\text{H}}=E_{\text{Janus-TMDC + H}}-E_{\text{Janus-TMDC}}-\frac{1}{2}E_{\text{H}},$ (16)
where $E_{\text{Janus-TMDC + H}}$ is the total energy of the heterostructure of Janus monolayer and TMDC material with an adsorbed hydrogen atom, $E_{\text{Janus-TMDC}}$ is the energy of the bare vdW layers, and $\frac{1}{2}E_{\text{H}}$ represents the energy of a hydrogen molecule in its standard state. Building upon this, the potential provided by photogenerated electrons for HER ($U_{e}$) can be calculated using the equation $U_{e}=V_{\text{CB}}-4.027$ for pH = 7 14, 63, where $U_{e}$ corresponds to the electron potential associated with the CBM, potentially matching the electron affinity $\chi(\text{H}_{2})$ under illuminated conditions. In a dark environment, where no photogenerated electrons are available, $U_{e}$ becomes zero, indicating the absence of a driving force for HER. Based on the calculations in the previous section, the $U_{e}$ values for $\text{WS}_{2}$-SMoSe, $\text{WSe}_{2}$-TeWSe, and $\text{WS}_{2}$-SMoSe under the illumination were found to be 0.10, 0.12, and 0.09 eV, respectively.
We further evaluated the HER activity at the reduction sites of $\text{WSe}_{2}$-SWSe, $\text{WSe}_{2}$-TeWSe, and again $\text{WS}_{2}$-SMoSe, with the corresponding $\Delta G_{\text{H}}$ values presented in Figure 5. The calculated free energy changes without an external potential are 2.10, 2.18, and 2.27 eV for these materials, respectively. These high positive values of $\Delta G_{\text{H}}$ indicate that spontaneous hydrogen adsorption does not occur easily, implying limited catalytic efficiency under standard conditions. Nevertheless, this behavior is similar to other well-known materials such as $\text{MoS}_{2}$, $\text{MoSe}_{2}$, $\text{WSe}_{2}$, and the Janus MoSSe materials 64, 65, 66. On the other hand, calculated free energy changes with an external potential are the $\Delta G_{\text{H}}$ values to 2.01, 2.06, and 2.17 eV for $\text{WSe}_{2}$-SWSe, $\text{WSe}_{2}$-TeWSe, and $\text{WS}_{2}$-SMoSe, respectively. These reduction values suggest that suitable irradiation can enhance the HER performance by decreasing the energy barrier for hydrogen adsorption. Although the $\Delta G_{\text{H}}$ values remain positive, indicating that the hydrogen evolution reaction is still non-spontaneous, the decreased values under light exposure or applied potential could potentially improve the catalytic performance by making hydrogen releasing more favorable.
3.6.2 Diffusion of H atom on the HER sites
To investigate H atom diffusion on the HER-active surfaces, we calculate the optimized migration pathways using a minimal energy approach 45. The pathways and the corresponding energy profiles are shown in Figure 6, where S1 $\rightarrow$ S2 represents H atom migration between adjacent hex-atomic-rings, and S2 $\rightarrow$ S3 denotes migration within a single hex-atomic-ring. The migration pathways are optimized by the nudged elastic band (NEB) method 45 with a convergence criterion of 0.05 eV/Å for NEB calculations with 13 intermediate images. For WS2-SMoSe, H migration on the SMoSe surface showed a lower energy barrier for inter-ring migration (0.199 eV) than intra-ring migration (0.216 eV). In the case of WSe2-TeWSe, the WSe2 surface exhibited easier migration between hex-atomic-rings (0.149 eV) than within a ring (0.286 eV). For WS2-SMoSe, the SMoSe surface favored intra-ring migration (0.200 eV) over inter-ring movement (0.224 eV), with this structure exhibiting a lower energy barrier compared to the other two. Compared to previously reported values 45, the migration energy barriers of these structures are lower than those of monolayer MoSSe, which were observed at 0.32 eV for inter-ring and 0.08 eV for intra-ring migration on Se-site. This reduction suggests that these WS2-SMoSe, WSe2-TeWSe, and WS2-SMoSe surfaces provide more favorable diffusion pathways, which could enhance HER activity on these materials.
Figure 6: Migration pathways and energy profiles of a Hydrogen atom on the reduction site: (a) Janus site (WSe2-SWSe and WS2-SMoSe) and (b) TMDC site (WSe2-TeWSe). S1 $\rightarrow$ S2 indicates migration between neighboring hex-atomic-rings, while S2 $\rightarrow$ S3 represents migration within a hex-atomic-ring.
4 Conclusions
The findings in this study underscore the potential of Janus-TMDC heterobilayers as highly efficient photocatalysts for STH conversion. By leveraging intrinsic electric fields generated through Janus asymmetry, we achieve enhanced charge separation and minimized electron-hole recombination in Z-scheme configurations such as WSe2-SWSe, WSe2-TeWSe, and WS2-SMoSe. These Z-scheme structures not only enable spatially separated sites for HER and OER but also provide favorable band alignment for water splitting, with WS2-SMoSe achieving the highest STH efficiency of 33.24%. In exploring the relationship between carrier mobility and photocatalytic capabilities, we use the Fröhlich interaction model to account for LO phonon scattering, revealing that mobility in Janus-TMDC heterobilayers is significantly affected by their intrinsic polarization. Additionally, HER analysis shows that external potential effectively lowers reaction barriers, enhancing HER efficiency under illumination. Hydrogen diffusion also indicates relatively low energy barriers for H atom migration on HER-active surfaces, suggesting efficient diffusion on the WSe2-SWSe, WSe2-TeWSe, and WS2-SMoSe surfaces. Overall, this work provides a comprehensive framework for selecting and designing the optimized Janus-TMDC heterobilayers for clean energy applications, presenting a promising pathway toward advancing efficient photocatalytic systems.
{acknowledgement}
This work was financially supported by Vietnam National University Ho Chi Minh City (NCM2024-50-01).
{suppinfo}
Supporting Information: Table of structural parameters and band Gap; table of Born charges, Phonon frequency, and dielectric properties; figure of HSE06-calculated band structure; figure of electrostatic potential along vacuum axis ($z$); figure of band alignments for water splitting at pH 0, 2, and 7; figure of PBE band structure for the effective mass calculation; figure of total energy-strain relationship in zigzag and armchair directions; figure of CBM and VBM energies under uniaxial strain; figure of Phonon dispersion and DOS.